### v1

# Zens, Díaz, Durante and Patacchini

Abstract
Criminal activity data are typically available via a three-way tensor encoding the reported frequencies of different crime categories across time and space. The challenges that arise in the design of interpretable, yet realistic, model-based representations of the complex dependencies within and across these three dimensions have led to an increasing adoption of black-box predictive strategies. Although this perspective has proved successful in producing accurate forecasts guiding targeted interventions, the lack of interpretable model-based characterizations of the dependence structures underlying criminal activity tensors prevents from inferring the cascading effects of these interventions across the different dimensions. We address this gap through the design of a low-rank bilinear autoregressive model which achieves comparable predictive performance to black-box strategies, while allowing interpretable inference on the dependence structures of criminal activity reports across crime categories, time, and space. This representation incorporates the time dimension via an autoregressive construction, accounting for spatial effects and dependencies among crime categories through a separable low-rank bilinear formulation. When applied to Chicago police reports, the proposed model showcases remarkable predictive performance and also reveals interpretable dependence structures unveiling fundamental crime dynamics. These results facilitate the design of more refined intervention policies informed by cascading effects of the policy itself.
keywords:
Bayesian statistics, criminal activity data, graphical models, matrix-variate time series
1 1. Introduction
The empirical analysis of criminal activities plays an important role in both theory development and public policy considerations. Theoretical models of crime indicate that criminal incidents do not occur randomly, but rather exhibit complex interdependencies across spatial areas, time, and crime categories (e.g., Brantingham and Brantingham, 2016; Weisburd, 2015). As a result, several methods have been proposed to characterize the intricate system of interdependencies within and across these three dimensions. Available solutions for addressing this endeavor range from interpretable model-based representations (e.g., Shoesmith, 2013; Law et al., 2014; Liu and Zhu, 2017; Quick et al., 2018; Clark and Dixon, 2018; Balocchi and Jensen, 2019; Vicente et al., 2023; Balocchi et al., 2023) to more sophisticated black-box predictive strategies (e.g., Guo and Wu, 2012; Chun et al., 2019; Walczak, 2021; Berk, 2021; Zhang et al., 2022; Liang et al., 2022; Mandalapu et al., 2023; Wu et al., 2024; Wang et al., 2025). Although statistical models allow, in principle, for interpretable inference on the complex dependence structures underlying the observed criminal activity tensors, the validity of the final conclusions is often undermined by the challenges that arise in the design of sufficiently-flexible, yet interpretable, realistic model-based representations for these dependence structures. As such, the existing statistical models often constrain the analysis by focusing on two dimensions of the three-way tensor, fixing the third at a selected crime category, a single geographic location, or a pre-specified time interval. Although such a perspective has led to important advancements, as discussed above, constraining the analysis only to two dimensions fails to characterize the full set of interdependencies underlying the criminal activity tensor, thereby limiting the policy impact of the resulting inferences. Addressing these issues would require extending the available flexible models for two dimensions to similarly-effective formulations for the entire three-way tensor. In fact, instead of exploring such a direction, most efforts in the recent years have been devoted to the development and implementation of black-box strategies within a predominantly prediction-oriented framework (see, e.g., Berk, 2021; Mandalapu et al., 2023). These strategies have proved successful in producing accurate forecasts guiding targeted intervention policies, such as hot spot policing (Weisburd and Telep, 2014; Braga et al., 2014), but generally lack a model-based construction, and therefore make it challenging to derive meaningful structural insights on the cascading, positive or negative, spillover effects of specific localized interventions (see, e.g., Weisburd et al., 2006; Blattman et al., 2021). Learning these effects is fundamental to devise improved intervention policies that not only rely on the forecasted criminal activities, but are also guided by the expected cascading effects of the planned policy itself.
Motivated by the above discussion and challenges, we design within Section 3 a model-based Bayesian representation of three-way criminal activity tensors that achieves a predictive performance comparable to routinely-used black-box strategies, while allowing for interpretable inference on the complex dependence structures within and across crime categories, spatial locations and time intervals. This formulation is guided by qualitative and quantitative theories of crime discussed in Section 2, and relies on a temporal autoregressive construction that incorporates spatial effects and graphical dependencies among different crime categories via a separable low-rank bilinear formulation inspired by the general class of multilinear tensor regression models in Hoff (2015). Besides reducing the dimensionality of the problem, this novel perspective to the analysis of crime activity tensors crucially unveils sparse averaged inter-temporal dependencies between crime categories within spatial locations as well as among spatial locations for each crime category. Such a separable structure is also considered for the noise component of the model, thereby further uncovering across-categories and across-locations immediate reactions to shocks, possibly arising from targeted hot spot interventions. As clarified in Section 4, inference under the proposed model is performed through a Bayesian approach leveraging a suitably-designed Gibbs sampling algorithm. The advantages of our formulation are illustrated within Section 5 in an application to a three-way criminal activity tensor obtained from the Chicago Police Department. Besides achieving a predictive performance comparable to routinely-implemented black-box strategies, the proposed model unveils yet-unexplored statistical interdependencies between crime categories and within locations over time that augment the set of information available to those policymakers seeking to understand and prevent complex patterns in criminal activities. As discussed in the concluding remarks within Section 6, our novel approach can be applied in general to criminal activity tensors from any urban environment for which similar data are available.
2 2. Qualitative and Quantitative Evidence on Criminal Activities
2.1 2.1. Determinants and mechanisms
Crime is mainly an urban phenomenon. Urban agglomerations typically exhibit higher crime rates compared to smaller towns and rural areas (e.g., Glaeser and Sacerdote, 1999). Although larger cities are often associated with higher economic development, and hence more attractive, this is not the only dimension that explains the distribution of crime across cities. In fact, crime reports are typically sorted unequally across different urban locations, with offenses often concentrated by type in relatively small areas. This spatial heterogeneity, coupled with specific temporal persistences, has led social scientists to focus on local determinants rather than individual drivers.
Three traditional criminology perspectives are useful for understanding the connections among time, space and crime categories. Crime opportunity theories, which focus on the factors shaping space-time crime patterns, particularly the clustering of criminal activities (e.g., Brantingham and Brantingham, 2017; Cohen and Felson, 1979); social control theories, where the structural dimensions of a neighborhood influence residents propensity to exert informal social control (Sampson and Groves, 1989; Shaw and McKay, 1942); and theories that simply highlight how local characteristics (e.g., medical marijuana dispensaries, parks, neighborhood stability and socioeconomic disparities) affect incentives for committing a specific crime within a given area (Contreras, 2017; Contreras and Hipp, 2020; Kubrin and Hipp, 2016). It is important to note that crime clusters are not driven by a single factor in isolation (Hipp and Williams, 2020). Rather, the interplay between opportunity, social controls, and local conditions causes urban areas to experience varying levels of temporal exposure to specific types of crimes and violence.
To illustrate the complexity of the determinants and mechanisms linking time, space, and crime categories it is possible to recall the routine activity theory. Originally developed by Cohen and Felson (1979), this perspective offers one of the clearest frameworks for explaining the relationship between opportunity and rational choice (Becker, 1968; Ehrlich, 1973). According to this theory, crime is unevenly distributed (Weisburd, 2015) and occurs when three key conditions converge in time and space, namely motivated offenders, suitable targets (valuable objects or potential victims), and the absence of effective interventions to prevent crime (Felson and Eckert, 2018). As a result, the location and persistence of crime within urban areas are not random but rather shaped by specific situational contexts, which influence how offenders perceive the costs and benefits of criminal activity (Becker, 1968). Crime has higher chance to occur when offenders access potential targets without interference, and since this convergence tends to vary by crime type, distinct spatio-temporal patterns emerge for different offenses. Hence, the routine activity approach offers a valuable framework to motivate and support the design of models capable of disentangling the complex system of interdependencies in the three-way criminal activity tensor.
Although it is challenging to separate time and space when analyzing urban crime patterns, the temporal dimension of crime determinants has been documented to play a crucial role in shaping its spatial distribution, alongside factors related to the offender behavior. Several criminogenic features of a geographic area tend to remain stable over time, contributing to the persistence of specific criminal activities (Johnson, 2008). For example, neighborhood characteristics that either promote or inhibit social control are likely to remain constant (Sampson and Groves, 1989). In addition, the success of previous offenses often motivates criminals to continue operating in the same area, capitalizing on the knowledge acquired about available targets and absence of countermeasures (Cohen and Felson, 1979). The literature on near repeat victimization further suggests that offenders are more likely to target locations where similar crimes have been successfully committed in the past (Johnson and Bowers, 2004; Townsley et al., 2003). However, not all crimes are equally susceptible to temporal persistence and spatial clustering (e.g., Pease et al., 1998; Johnson and Bowers, 2004; Ratcliffe and Rengert, 2008; Johnson et al., 2009; Haberman and Ratcliffe, 2012; Amemiya et al., 2020; Chainey, 2021; Chainey and Muggah, 2022). In fact, as discussed in these studies, the clustering of urban crime in time and space is shaped by patterns that are heterogeneous across crime categories. In specific urban areas, certain types of offenses are more prevalent than others, leading to a localized concentration of particular crimes (e.g., property crimes versus violent crimes). These patterns are shaped by the spatial distribution of suitable targets and by the presence of deterrence interventions (Cohen and Felson, 1979), which can favor the occurrence of some crimes over others. For instance, the covid-19 pandemic highlighted the role of crime opportunity in explaining the global reduction in street crime (Nivette et al., 2021; Trajtenberg et al., 2024), while at the same time, the literature documented increases in domestic violence and cybercrime (Piquero et al., 2021; Ravindran and Shah, 2023). Empirical evidence suggests that offenders may have increased the use of violence to compensate for the reduced opportunities brought on by the pandemic, thereby reshaping the spatial distribution and clustering of violent and non-violent crimes (Díaz et al., 2022). Another example of crime-type clustering heterogeneity results from the strategic decision-making of organized crime groups when selecting or relocating to new territories (Varese, 2011). These groups typically target urban communities that offer favorable conditions for the development of illegal markets, display socio-economic deprivation, and demonstrate a higher degree of homophily (Campana and Meneghini, 2024; Katz and Schnebly, 2011). For instance, gang members are more likely to settle in communities that share similar social and demographic traits. This dynamic helps in explaining why activities such as production, trafficking, and governance (Breuer and Varese, 2023; Campana and Varese, 2018), along with related criminal behaviors, are more prevalent in certain urban areas than in others.
2.2 2.2. Policy implications
The primary policy implication of the geographic concentration and temporal persistence of crime mainly involves the adoption of targeted deterrence strategies, with a particular emphasis on policing. As most crimes are concentrated in small geographic areas (Telep and Weisburd, 2018; Weisburd et al., 2004), law enforcement approaches must be tailored to this reality. One of the most popular strategies is the hot spots approach (e.g., Weisburd and Telep, 2014; Braga et al., 2014), where police resource allocation focuses on geographic areas with a higher number of crimes than the city’s average or where people are at a higher risk of victimization. In this way, police resources could be allocated to just 5% of its block segments to address more than 50% of the total reported incidents it receives (Weisburd, 2015). This approach is mostly guided by predictive techniques that forecast crime in both time and space (Berk, 2021), while the response may involve strategies such as increasing foot patrols, engaging with the local stakeholders, or implementing problem-oriented policing (Braga and Bond, 2008; Taylor et al., 2011; Weisburd et al., 2022).
The main issue with the above perspective is that an over-reliance on purely-predictive methods fails to inform on the cascading effects of hot spot interventions, which include the risk of crime and disorder displacement. Critics of intensive hot spot policing argue that such strategies tend to displace crime rather than effectively reduce it. The literature is inconclusive and cannot rule out positive or negative spillover effects. Braga et al.’s (2019) review shows that studies usually do not detect noticeable negative spillover to other areas. Several analyses even identify benefits that spread from hot spots to adjacent areas (e.g., Weisburd et al., 2006). Both incapacitation and deterrence can explain these results. If the arrested offenders also operated in neighboring areas, the offense is expected to decrease in the hot spots and surrounding areas. On the other hand, some studies document crime displacement when the police strategy focuses on specific areas. For example, a study by Blattman et al. (2017) finds that any crime deterrence in the target area is more than offset by increased crime on the streets within the following 250 meters. In addition, such an analysis unveils heterogeneous effects by crime category. While violent crimes decrease, property crimes increase. The authors conclude that this type of policing pushes property crime around the corner. Similarly, Banerjee et al. (2019) use randomized drunk driving checkpoints to find learning and strategic responses that explain the displacement of crime to other areas.
The lack of a consensus in the aforementioned research studies underscores the pressing need for a more comprehensive understanding of how various crime categories manifest, and interact, in different urban areas across time. Even though crime clustering in time and space is a known phenomenon, an unanimous consensus remains elusive regarding the specific interactions among criminal activities in time, space and across different crime categories. Consequently, an ongoing debate surrounds the tangible benefits of specific intervention policies, particularly those with a pronounced local emphasis (e.g., the hot spots approach) based on black-box predictive methods. In fact, such methods are not designed to provide a model-based and interpretable understanding of complex dependencies in three-way criminal activity tensors. In Section 3, we cover such a gap through a statistical model that can disentangle these dependencies, while preserving a predictive performance competitive with routine-use black-box methods. As such, the proposed perspective can open the avenues for refining the design and evaluation of law enforcement interventions by combining accurate forecasts with a detailed quantitative assessment of the intended or unintended spillover effects that these interventions may trigger. To the best of our knowledge, this important direction has received limited attention in the literature.
2.3 2.3. Existing statistical approaches
As discussed within Section 1, available model-based solutions for the study of criminal activity data (see, e.g., Shoesmith, 2013; Law et al., 2014; Liu and Zhu, 2017; Quick et al., 2018; Clark and Dixon, 2018; Balocchi and Jensen, 2019; Vicente et al., 2023; Balocchi et al., 2023) mostly constrain the analysis to two selected dimensions of the observed three-way tensor. Although such a perspective fails to characterize and learn the full range of dependencies across time, space and crime categories, the available contributions along these directions highlight several fundamental aspects that are the basis of the statistical model we design in Section 3.
More specifically, in line with the discussion in Sections 2.12.2, the design of careful model-based representations for the complex dependencies across time and space is crucial in analyzing criminal activity data. For example, Shoesmith (2013) demonstrates the importance of temporal lags and spatio-temporal interactions when modeling criminal activities. Similarly, Balocchi and Jensen (2019) and Balocchi et al. (2023) show how careful borrowing of information among temporal crime trends across space improves modeling outcomes. Vicente et al. (2023) employ, instead, a penalized spline expansion across locations and times, further allowing for joint modeling of different crime categories. Although such a formulation does not explicitly incorporate graphical dependencies among these different crime categories, it provides relevant support to the importance of devising joint formulations for the entire three-way criminal activity tensor. This support can be found also in the static spatial models by, e.g., Liu and Zhu (2017) and Quick et al. (2018), which highlight how a multivariate perspective focusing on multiple crime categories can unveil structural patterns that a univariate model is unable to detect. Although black-box strategies (see, e.g., Berk, 2021; Mandalapu et al., 2023) can jointly leverage this complex system of dependencies within the observed three-way criminal activity tensor to achieve a remarkable predictive accuracy, the design of a statistical model that can disentangle and infer these dependencies, without sacrificing forecasting performance, is substantially more challenging. Although such a perspective has received limited attention, as discussed in Section 2.2, advancements in this direction are fundamental to devise strategies based not only on accurate forecasts, but also on a careful assessment of the cascading effect of the adopted intervention strategy. The relevance of this endeavor has motivated some efforts towards providing structural interpretation to black-box algorithms in crime analysis, but these interpretations are mostly based on ex-post variable importance measures in the case where external covariates are related to crime activity data under some sort of regression structure (e.g., Zhang et al., 2022).
In Section 3, we address the above issues via a Bayesian model-based formulation that balances parsimony, interpretability and predictive accuracy by integrating ideas from matrix autoregressive processes, bilinear constructions, low-rank factorizations and graphical models.
3 3. Low-Rank Bilinear Autoregressive Model for Criminal Activities
Following similar practice in the analysis of criminal activity data ( e.g., Shoesmith, 2013; Balocchi and Jensen, 2019), we model the counts $c_{kqt}\in\mathbbm{N}$ of the reported crimes for category $k=1,\dots,K$, within location $q=1,\dots,Q$ at time $t=1,\dots,T$ under the $\log(1+c_{kqt})$ transformation. Let $y_{kqt}$ be the centered target variable obtained by removing from $\log(1+c_{kqt})$ its average over $t$ (keeping $k$ and $q$ fixed), and denote with $\bf{Y}$ the $K\times Q\times T$ criminal activity tensor having entries $y_{kqt}$, for each $k=1,\dots,K$, $q=1,\dots,Q$ and $t=1,\dots,T$. In Section 3.1, we design a statistical model for $\bf{Y}$ which characterizes the temporal dependence among its $K\times Q$ time slices ${\bf Y}_{t}$, $t=1,\dots,T$, via a matrix autoregressive construction that effectively separates space and crime-category effects through a parsimonious, yet interpretable, low-rank bilinear formulation.
The above model, combined with a separable covariance structure for the error term, allows to infer spatial dependence, interlinkages between the different crime categories and persistence over time, without undermining predictive accuracy. Inference for the proposed model proceeds under a Bayesian framework outlined in detail in Section 4.
3.1 3.1. Model formulation
To devise a model that captures spatial, temporal and crime-category effects within the three-way criminal activity tensor, while facilitating the design of forecasting strategies, a simple building-block option is to rely on a full autoregressive construction for the matrix-valued sequence ${\bf Y}_{t}$, $t=1,\dots,T$. More specifically, let ${\bf y}_{t}=[y_{11t},\dots,y_{kqt},\dots y_{KQt}]^{\intercal}$ denote the vectorized version of the matrix ${\bf Y}_{t}$, for each $t=1,\dots,T$, then it is natural to start by assuming ${\bf y}_{t}=\bm{\Theta}{\bf y}_{t-1}+\bm{\varepsilon}_{t}$, with $\bm{\varepsilon}_{t}\sim\mbox{N}_{KQ}({\bf 0},\bm{\Sigma})$, independently for every $t=1,\dots,T$. In this standard vector autoregressive formulation the lagged values within ${\bf y}_{t-1}$ affect the current ones in ${\bf y}_{t}$ via a full and unstructured $KQ\times KQ$ coefficient matrix $\bm{\Theta}$, which captures the different inter-temporal effects that every combination of location and crime category exerts on each entry in ${\bf y}_{t}$, whereas the $KQ\times KQ$ error covariance matrix $\bm{\Sigma}$ encodes the instantaneous dependencies in reaction to possible shocks.
Although unstructured vector autoregressive constructions are not novel in the statistical analysis of criminal activity data (e.g., Funk and Kugler, 2003; Shoesmith, 2013), none of the available solutions is designed to study the time changes of criminal activities jointly across both space and crime categories. In fact, current solutions focus only on one of the two dimensions. As discussed in Section 2, this choice is practically and conceptually suboptimal, since there are nuanced interactions among types of crime and spatial locations across time. Accounting explicitly for both, as in the above model, facilitates borrowing of information to improve forecasts and opens the avenues to provide interpretable inference on these interactions. To accomplish this goal it is however important to consider a structured representation for both the coefficient matrix $\bm{\Theta}$ and the error covariance matrix $\bm{\Sigma}$ in order to disentangle spatial and crime-category dependencies, while avoiding overparameterized formulations with potential overfitting issues.
Motivated by the above considerations, we rely on a separable construction that assumes $\bm{\Theta}={\bf B}\otimes{\bf A}$ and $\bm{\Sigma}=\bm{\Sigma_{B}}\otimes\bm{\Sigma_{A}}$, where $\otimes$ denotes the Kronecker product. In this factorization, ${\bf A}$ and ${\bf B}$ are $K\times K$ and $Q\times Q$ matrices that capture the average inter-temporal dependence between the $K$ crime categories within spatial locations, and among these $Q$ spatial locations within each crime category, respectively. Similarly, $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$ correspond to the $K\times K$ and $Q\times Q$ covariance matrices encoding the across-categories and across-locations instantaneous dependencies in reaction to shocks, respectively. In such a way, it is possible to disentangle spatial and crime-category dependencies, thereby enhancing interpretability, while preserving flexibility by letting such dependencies interact in the definition of $\bm{\Theta}$ and $\bm{\Sigma}$. Separable constructions have been successfully employed in black-box methods (e.g., Wang et al., 2025) to improve predictive accuracy, but the potential of these perspectives to allow for interpretable model-based inference on the complex dependencies within the criminal activity tensor have not yet been explored.
Leveraging properties of the Kronecker product and of matrix-variate normal (MN) distributions, the above separable autoregressive construction admits the following bilinear representation
${\bf Y}_{t}={\bf A}{\bf Y}_{t-1}{\bf B}^{\intercal}+{\bf E}_{t},\qquad{\bf E}_{t}\stackrel{{\scriptstyle\mbox{\footnotesize i.i.d.}}}{{\sim}}\mbox{MN}_{K,Q}({\bf 0},\bm{\Sigma}_{A},\bm{\Sigma}_{B}),$ (1)
for every $t=1,\ldots,T$. This formulation specializes the general class of multilinear tensor regression models in Hoff (2015) to our specific setting by considering a matrix-valued outcome related to a matrix-variate input via a temporal autoregressive structure; see also Billio et al. (2023) and the references therein for more general multi-way tensor autoregression constructions with a main focus on econometric applications. Crucially, compared to a fully-unstructured vector autoregressive formulation, the model in (1) reduces the number of parameters from $(KQ)^{2}+(KQ)(KQ+1)/2$ to $K^{2}+Q^{2}+K(K+1)/2+Q(Q+1)/2$. In our application to Chicago criminal activity data with $K=28$ crime categories and $Q=22$ spatial locations, this implies a reduction from $569{,}492$ parameters to $1{,}927$ (about three hundred times lower than those within the unstructured model). Interestingly, as illustrated in Section 5, such a remarkable reduction in complexity still yields a predictive performance competitive with routine-use black-box methods, meaning that model (1) properly accounts for a rich set of dynamics within and between crime categories, locations and times.
Although the above formulation achieves a sensible balance between predictive potential and interpretable inference on the dependencies in the criminal activity tensor, in practice, unrestricted estimation of the $K\times K$ and $Q\times Q$ matrices ${\bf A}$ and ${\bf B}$ may still result in a overly-parametrized model. In addition, as discussed in Section 2, it is reasonable to expect that both ${\bf A}$ and ${\bf B}$ would admit lower-dimensional representations due the sparse and structured inter-temporal dependencies associated with both crime categories and spatial locations. To account for this information and further reduce complexity, we combine model (1) with a low-rank factorization for ${\bf A}$ and ${\bf B}$ defined as
${\bf A}={\bf L}_{A}{\bf Z}_{A},\qquad\qquad{\bf B}={\bf L}_{B}{\bf Z}_{B},$ (2)
where ${\bf L}_{A}$ and ${\bf L}_{B}$ have dimension $K\times R_{A}$ and $Q\times R_{B}$, respectively, with $R_{A}\leq K$ and $R_{B}\leq Q$. Conversely, the matrices ${\bf Z}_{A}$ and ${\bf Z}_{B}$ are of dimension $R_{A}\times K$ and $R_{B}\times Q$. In this way, the fixed integers $R_{A}$ and $R_{B}$ determine the rank of the coefficient matrices ${\bf A}$ and ${\bf B}$, thereby further reducing the number of free parameters in both ${\bf A}$ and ${\bf B}$. This allows for additional reductions in complexity compared to a full-rank matrix-autoregressive model. In the motivating application, $R_{A}$ and $R_{B}$ are specified based on the results of an out-of-sample predictive exercise. See Reinsel et al. (2022) and Tsay (2024) for related factorizations in reduced-rank multivariate regression models and for recent developments in matrix-variate autoregressive settings.
Despite the potential of the above ideas for learning structure in three-way criminal activity tensors, to our knowledge formulations of this type have been overlooked within criminology settings. To cover such a gap we employ the model in (1)–(2) and perform inference on its parameters under a Bayesian perspective (see Section 4). Besides introducing additional shrinkage effects through prior distributions, this choice provides a fully probabilistic framework enabling formal uncertainty quantification, also in forecasting, and possible inclusion of expert information along with nuanced hierarchical structures for more sophisticated crime modeling exercises.
4 4. Bayesian Inference and Computation
4.1 4.1. Prior specification
To perform Bayesian inference under model (1)–(2) it is necessary to elicit prior distributions for its parameters, namely $\bm{\Sigma}_{A}$, $\bm{\Sigma}_{B}$, ${\bf L}_{A}$, ${\bf L}_{B}$, ${\bf Z}_{A}$ and ${\bf Z}_{B}$. These priors are specified to facilitate tractable posterior inference, while introducing mild shrinkage effects that, however, do not overwhelm the information provided by the observed criminal activity tensor through the likelihood of model (1)–(2).
Consistent with the above goals, we follow routine-practice in Bayesian modeling of covariance matrices (e.g., Gelman et al., 2013) by relying on inverse-Wishart priors for $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$; namely
$\bm{\Sigma}_{A}\sim\mbox{inv-Wishart}(4,2\cdot{\bf I}_{K}),\qquad\quad\bm{\Sigma}_{B}\sim\mbox{inv-Wishart}(4,2\cdot{\bf I}_{Q}).$
Combined with the Gaussian likelihood induced by model (1)–(2), these priors admit conditional-conjugacy properties that facilitate Bayesian inference on the instantaneous dependence structures among the $K$ crime categories and between the $Q$ spatial locations. Notice that, although the hyperparameters of the above priors can be possibly set at different values, in the application in Section 5 we found these choices effective both in inference and in out-of-sample prediction.
As for the priors on ${\bf Z}_{A}$ and ${\bf Z}_{B}$, we adapt to our specific low-rank bilinear autoregressive setting the elicitation suggested for the general class of multilinear tensor regression models in Hoff (2015). To this end, we let
${\bf Z}_{A}\sim\mbox{MN}_{R_{A},K}({\bf 0},{\bf I}_{R_{A}},10\cdot{\bf I}_{K}),\qquad\quad{\bf Z}_{B}\sim\mbox{MN}_{R_{B},Q}({\bf 0},{\bf I}_{R_{B}},10\cdot{\bf I}_{Q}),$
thereby inducing via the low-rank construction in (2) weakly-diffuse and conditionally-conjugate priors on the separable inter-temporal effects in ${\bf A}$ and ${\bf B}$. Under (2), such effects are also controlled by ${\bf L}_{A}$ and ${\bf L}_{B}$ for which we assume flat priors combined with a processing step within the proposed Gibbs sampler that enforces the constraints ${\bf L}_{A}^{\intercal}{\bf L}_{A}={\bf I}_{R_{A}}$ and ${\bf L}_{B}^{\intercal}{\bf L}_{B}={\bf I}_{R_{B}}$, respectively. This choice implicitly fixes the scale of the elements within ${\bf L}_{A}$ and ${\bf L}_{B}$ thereby addressing rotational and scaling identification issues that may arise from the factorization in (2). While these identifiability problems only affect ${\bf L}_{A}$, ${\bf Z}_{A}$ and ${\bf L}_{B}$, ${\bf Z}_{B}$, and not the main objects of inference, i.e., ${\bf A}$ and ${\bf B}$, considering processing steps that address these potential issues improves mixing and convergence of the Gibbs sampler we design in Section 4.2 for posterior inference.
4.2 4.2 Posterior computation via Gibbs sampler
The prior distributions in Section 4.1 combined with the likelihood of model (1)–(2) yield a posterior for the parameters $\bm{\Sigma}_{A}$, $\bm{\Sigma}_{B}$, ${\bf L}_{A}$, ${\bf L}_{B}$, ${\bf Z}_{A}$ and ${\bf Z}_{B}$ which is not available in closed form. Nonetheless, as clarified below, tractable full conditionals for each of these parameters given the others and the observed data are available, thereby allowing for Monte Carlo inference leveraging samples from the posterior of these parameters obtained via a Gibbs sampling algorithm.
To derive in detail the different steps of such a sampling algorithm, let ${{\bf Y}}_{(A)}=[{\bf Y}_{1},\ldots,{\bf Y}_{T}]$ and ${{\bf Y}}_{(B)}=[{\bf Y}^{\intercal}_{1},\ldots,{\bf Y}^{\intercal}_{T}]$. Moreover, denote with ${{\bf X}}_{(A)}$ and ${{\bf X}}_{(B)}$ the lagged version of ${{\bf Y}}_{(A)}$ and ${{\bf Y}}_{(B)}$ respectively. Then, under these settings, and conditionally on ${\bf B}$ and $\bm{\Sigma}_{B}$, model (1) can be expressed jointly over the whole time window as
$\tilde{{\bf Y}}_{(A)}={\bf A}{\tilde{\bf X}}_{(A)}+\tilde{{\bf E}}_{(A)},\quad\quad\tilde{{\bf E}}_{(A)}\sim\mbox{MN}_{K,TQ}({\bf 0},\bm{\Sigma}_{A},{\bf I}_{TQ}),$ (3)
with $\tilde{{\bf Y}}_{(A)}={{\bf Y}}_{(A)}\bm{\Sigma}_{B}^{-1/2}$ and ${\tilde{\bf X}}_{(A)}={{\bf X}}_{(A)}{\bf B}^{\intercal}\bm{\Sigma}_{B}^{-1/2}$, where $\bm{\Sigma}_{B}^{-1/2}$ is the Cholesky decomposition of the inverse of $\bm{\Sigma}_{B}$. We refer to representation (3) as $A$-matricization of (1). Replacing ${\bf A}$ with its low-rank factorization in (2) and leveraging standard properties of vectorizations, such a representation can be also expressed, in vector form, as
$\text{vec}(\tilde{{\bf Y}}_{(A)})=({\tilde{\bf X}}_{(A)}^{\intercal}\otimes{\bf L}_{A})\text{vec}({\bf Z}_{A})+\text{vec}(\tilde{{\bf E}}_{(A)}),\quad\text{vec}(\tilde{{\bf E}}_{(A)})\sim\mbox{N}_{KTQ}({\bf 0},{\bf I}_{TQ}\otimes\bm{\Sigma}_{A}),$ (4)
or, equivalently, via
$\text{vec}(\tilde{{\bf Y}}_{(A)})=({\tilde{\bf X}}_{(A)}^{\intercal}{\bf Z}_{A}^{\intercal}\otimes{\bf I}_{K})\text{vec}({\bf L}_{A})+\text{vec}(\tilde{{\bf E}}_{(A)}),\quad\text{vec}(\tilde{{\bf E}}_{(A)})\sim\mbox{N}_{KTQ}({\bf 0},{\bf I}_{TQ}\otimes\bm{\Sigma}_{A}),$ (5)
thereby recasting model (1)–(2) as a Gaussian linear regression with coefficients $\text{vec}({\bf Z}_{A})$ (see (4)) or $\text{vec}({\bf L}_{A})$ (see (5)), respectively, conditionally on the other parameters. Combined with the priors for $\text{vec}({\bf Z}_{A})$ and $\text{vec}({\bf L}_{A})$ presented in Section 4.1, this formulation yields Gaussian full conditionals for both vectors. More specifically, $(\text{vec}({\bf Z}_{A})\mid-)\sim\mbox{N}_{R_{A}K}(\bm{\mu}_{{\bf Z}_{A}},\bm{\Sigma}_{{\bf Z}_{A}})$ with
$\begin{split}\bm{\Sigma}_{{\bf Z}_{A}}&=[10^{-1}{\bf I}_{R_{A}K}+({\tilde{\bf X}}_{(A)}^{\intercal}\otimes{\bf L}_{A})^{\intercal}({\bf I}_{TQ}\otimes\bm{\Sigma}_{A})^{-1}({\tilde{\bf X}}_{(A)}^{\intercal}\otimes{\bf L}_{A})]^{-1}\\ &=[10^{-1}{\bf I}_{R_{A}K}+({\tilde{\bf X}}_{(A)}{\tilde{\bf X}}_{(A)}^{\intercal}\otimes{\bf L}_{A}^{\intercal}\bm{\Sigma}_{A}^{-1}{\bf L}_{A})]^{-1},\\ \bm{\mu}_{{\bf Z}_{A}}&=\bm{\Sigma}_{{\bf Z}_{A}}({\tilde{\bf X}}_{(A)}^{\intercal}\otimes{\bf L}_{A})^{\intercal}({\bf I}_{TQ}\otimes\bm{\Sigma}_{A})^{-1}\text{vec}(\tilde{{\bf Y}}_{(A)})=\bm{\Sigma}_{{\bf Z}_{A}}\text{vec}({\bf L}_{A}^{\intercal}\bm{\Sigma}_{A}^{-1}\tilde{{\bf Y}}_{(A)}{\tilde{\bf X}}_{(A)}^{\intercal}).\end{split}$ (6)
Similarly, $(\text{vec}({\bf L}_{A})\mid-)\sim\mbox{N}_{R_{A}K}(\bm{\mu}_{{\bf L}_{A}},\bm{\Sigma}_{{\bf L}_{A}})$ with $\bm{\mu}_{{\bf L}_{A}}$ and $\bm{\Sigma}_{{\bf L}_{A}}$ defined as
$\begin{split}\bm{\Sigma}_{{\bf L}_{A}}&=[({\tilde{\bf X}}_{(A)}^{\intercal}{\bf Z}_{A}^{\intercal}\otimes{\bf I}_{K})^{\intercal}({\bf I}_{TQ}\otimes\bm{\Sigma}_{A})^{-1}({\tilde{\bf X}}_{(A)}^{\intercal}{\bf Z}_{A}^{\intercal}\otimes\bf{I}_{K})]^{-1}\\ &=[({\bf Z}_{(A)}{\tilde{\bf X}}_{(A)}{\tilde{\bf X}}_{(A)}^{\intercal}{\bf Z}_{(A)}^{\intercal}\otimes\bm{\Sigma}_{A}^{-1})]^{-1},\\ \bm{\mu}_{{\bf L}_{A}}&=\bm{\Sigma}_{{\bf L}_{A}}({\tilde{\bf X}}_{(A)}^{\intercal}{\bf Z}_{A}^{\intercal}\otimes{\bf I}_{K})^{\intercal}({\bf I}_{TQ}\otimes\bm{\Sigma}_{A})^{-1}\text{vec}(\tilde{{\bf Y}}_{(A)})=\bm{\Sigma}_{{\bf L}_{A}}\text{vec}(\bm{\Sigma}_{A}^{-1}\tilde{{\bf Y}}_{(A)}{\tilde{\bf X}}_{(A)}^{\intercal}{\bf Z}_{A}^{\intercal}).\end{split}$ (7)
Leveraging the above samples for $\text{vec}({\bf Z}_{A})$ and $\text{vec}({\bf L}_{A})$ it is possible to obtain directly a draw from the full-conditional distribution of the matrix ${\bf A}$ by computing ${\bf A}={\bf L}_{A}{\bf Z}_{A}$ at these samples. Conditioned on ${\bf A}$, it is then possible to leverage the conditional conjugacy of the inverse-Wishart prior for $\bm{\Sigma}_{A}$ under model (3), to obtain
$(\bm{\Sigma}_{A}\mid-)\sim\mbox{inv-Wishart}(4+TQ,2\cdot{\bf I}_{K}+(\tilde{{\bf Y}}_{(A)}-{\bf A}{\tilde{\bf X}}_{(A)})(\tilde{{\bf Y}}_{(A)}-{\bf A}{\tilde{\bf X}}_{(A)})^{\intercal}).$ (8)
A similar reasoning and derivations can be also considered to obtain closed-form and tractable full conditionals for ${\bf Z}_{B}$, ${\bf L}_{B}$ and $\bm{\Sigma}_{B}$. More specifically, given ${\bf A}$ and $\bm{\Sigma}_{A}$, it is possible to obtain the following $B$-matricization of (1)
$\tilde{{\bf Y}}_{(B)}={\bf B}{\tilde{\bf X}}_{(B)}+\tilde{{\bf E}}_{(B)}\quad\quad\tilde{{\bf E}}_{(B)}\sim\mbox{MN}_{Q,TK}({\bf 0},\bm{\Sigma}_{B},{\bf I}_{TK})$ (9)
where $\tilde{{\bf Y}}_{(B)}={{\bf Y}}_{(B)}\bm{\Sigma}_{A}^{-1/2}$ and ${\tilde{\bf X}}_{(B)}={{\bf X}}_{(B)}{\bf A}^{\intercal}\bm{\Sigma}_{A}^{-1/2}$, with $\bm{\Sigma}_{A}^{-1/2}$ denoting the Cholesky decomposition of the inverse of $\bm{\Sigma}_{A}$. Hence, as in (4)–(5), also in this case it is possible to express (9) in the form of a conditionally-Gaussian linear regression
$\text{vec}(\tilde{{\bf Y}}_{(B)})=({\tilde{\bf X}}_{(B)}^{\intercal}\otimes{\bf L}_{B})\text{vec}({\bf Z}_{B})+\text{vec}(\tilde{{\bf E}}_{(B)}),\quad\text{vec}(\tilde{{\bf E}}_{(B)})\sim\mbox{N}_{QTK}({\bf 0},{\bf I}_{TK}\otimes\bm{\Sigma}_{B}),$ (10)
with coefficients $\text{vec}({\bf Z}_{B})$, or via the alternative formulation
$\text{vec}(\tilde{{\bf Y}}_{(B)})=({\tilde{\bf X}}_{(B)}^{\intercal}{\bf Z}_{B}^{\intercal}\otimes{\bf I}_{Q})\text{vec}({\bf L}_{B})+\text{vec}(\tilde{{\bf E}}_{(B)}),\quad\text{vec}(\tilde{{\bf E}}_{(B)})\sim\mbox{N}_{QTK}({\bf 0},{\bf I}_{TK}\otimes\bm{\Sigma}_{B}),$ (11)
parameterized by $\text{vec}({\bf L}_{B})$. Combining (10) with the conditionally-conjugate Gaussian prior for ${\bf Z}_{B}$, implies that $(\text{vec}({\bf Z}_{B})\mid-)\sim\mbox{N}_{R_{B}Q}(\bm{\mu}_{{\bf Z}_{B}},\bm{\Sigma}_{{\bf Z}_{B}})$ with
$\begin{split}\bm{\Sigma}_{{\bf Z}_{B}}&=[10^{-1}{\bf I}_{R_{B}Q}+({\tilde{\bf X}}_{(B)}^{\intercal}\otimes{\bf L}_{B})^{\intercal}({\bf I}_{TK}\otimes\bm{\Sigma}_{B})^{-1}({\tilde{\bf X}}_{(B)}^{\intercal}\otimes{\bf L}_{B})]^{-1}\\ &=[10^{-1}{\bf I}_{R_{B}Q}+({\tilde{\bf X}}_{(B)}{\tilde{\bf X}}_{(B)}^{\intercal}\otimes{\bf L}_{B}^{\intercal}\bm{\Sigma}_{B}^{-1}{\bf L}_{B})]^{-1},\\ \bm{\mu}_{{\bf Z}_{B}}&=\bm{\Sigma}_{{\bf Z}_{B}}({\tilde{\bf X}}_{(B)}^{\intercal}\otimes{\bf L}_{B})^{\intercal}({\bf I}_{TK}\otimes\bm{\Sigma}_{B})^{-1}\text{vec}(\tilde{{\bf Y}}_{(B)})=\bm{\Sigma}_{{\bf Z}_{B}}\text{vec}({\bf L}_{B}^{\intercal}\bm{\Sigma}_{B}^{-1}\tilde{{\bf Y}}_{(B)}{\tilde{\bf X}}_{(B)}^{\intercal}),\end{split}$ (12)
while, (11) leads to $(\text{vec}({\bf L}_{B})\mid-)\sim\mbox{N}_{R_{B}Q}(\bm{\mu}_{{\bf L}_{B}},\bm{\Sigma}_{{\bf L}_{B}})$, where $\bm{\mu}_{{\bf L}_{B}}$ and $\bm{\Sigma}_{{\bf L}_{B}}$ are obtained as
$\begin{split}\bm{\Sigma}_{{\bf L}_{B}}&=[({\tilde{\bf X}}_{(B)}^{\intercal}{\bf Z}_{B}^{\intercal}\otimes{\bf I}_{Q})^{\intercal}({\bf I}_{TK}\otimes\bm{\Sigma}_{B})^{-1}({\tilde{\bf X}}_{(B)}^{\intercal}{\bf Z}_{B}^{\intercal}\otimes\bf{I}_{Q})]^{-1}\\ &=[({\bf Z}_{(B)}{\tilde{\bf X}}_{(B)}{\tilde{\bf X}}_{(B)}^{\intercal}{\bf Z}_{(B)}^{\intercal}\otimes\bm{\Sigma}_{B}^{-1})]^{-1},\\ \bm{\mu}_{{\bf L}_{B}}&=\bm{\Sigma}_{{\bf L}_{B}}({\tilde{\bf X}}_{(B)}^{\intercal}{\bf Z}_{B}^{\intercal}\otimes{\bf I}_{Q})^{\intercal}({\bf I}_{TK}\otimes\bm{\Sigma}_{B})^{-1}\text{vec}(\tilde{{\bf Y}}_{(B)})=\bm{\Sigma}_{{\bf L}_{B}}\text{vec}(\bm{\Sigma}_{B}^{-1}\tilde{{\bf Y}}_{(B)}{\tilde{\bf X}}_{(B)}^{\intercal}{\bf Z}_{B}^{\intercal}).\end{split}$ (13)
Finally, similarly to (8), the full conditional for $\bm{\Sigma}_{B}$ given ${\bf B}={\bf L}_{B}{\bf Z}_{B}$ is
$(\bm{\Sigma}_{B}\mid-)\sim\mbox{inv-Wishart}(4+TK,2\cdot{\bf I}_{Q}+(\tilde{{\bf Y}}_{(B)}-{\bf B}{\tilde{\bf X}}_{(B)})(\tilde{{\bf Y}}_{(B)}-{\bf B}{\tilde{\bf X}}_{(B)})^{\intercal}).$ (14)
Algorithm 1 Gibbs sampler for model (1)–(2)
for $s=1,\ldots,\mbox{S}$ do
     1. sample $\text{vec}({\bf Z}_{A})$ from the Gaussian full-conditional with parameters as in (6)
     2. sample $\text{vec}({\bf L}_{A})$ from the Gaussian full-conditional with parameters as in (7)
     3. transform the sampled ${\bf Z}_{A}$ and ${\bf L}_{A}$ to enforce the constraint discussed in Section 4.1
     4. sample $\bm{\Sigma}_{A}$ from the inverse Wishart full-conditional in (8)
     5. sample $\text{vec}({\bf Z}_{B})$ from the Gaussian full-conditional with parameters as in (12)
     6. sample $\text{vec}({\bf L}_{B})$ from the Gaussian full-conditional with parameters as in (13)
     7. transform the sampled ${\bf Z}_{B}$ and ${\bf L}_{B}$ to enforce the constraint discussed in Section 4.1
     8. sample $\bm{\Sigma}_{B}$ from the inverse Wishart full-conditional in (14)
     9. scale $\bm{\Sigma}_{A}$, $\bm{\Sigma}_{B}$, ${\bf A}=\mathbf{L}_{A}\mathbf{Z}_{A}$ and ${\bf B}=\mathbf{L}_{B}\mathbf{Z}_{B}$ to ensure identifiability
end for
Algorithm 1 summarizes the steps of the Gibbs sampler resulting from the above derivations. Notice that, due to the Kronecker structure of the matrix autoregressive model and the low-rank factorization of the coefficient matrices, some identifiability issues may arise. First, the matrix products $\mathbf{L}_{A}\mathbf{Z}_{A}$ and $\mathbf{L}_{B}\mathbf{Z}_{B}$ are only identifiable up to rotation, sign, scale and ordering of the respective rows and columns. Enforcing the constraints ${\bf L}_{A}^{\intercal}{\bf L}_{A}={\bf I}_{R_{A}}$ and ${\bf L}_{B}^{\intercal}{\bf L}_{B}={\bf I}_{R_{B}}$ discussed in Section 4.1 ensures scale and rotational identification, and increases numerical stability of the algorithm. Such constraints are included within the Gibbs sampler via an intermediate processing step that first computes the QR decomposition of both ${\bf L}_{A}$ and ${\bf L}_{B}$, i.e., ${\bf L}_{A}={\bf Q}_{L_{A}}{\bf R}_{L_{A}}$ and ${\bf L}_{B}={\bf Q}_{L_{B}}{\bf R}_{L_{B}}$, and then redefines ${\bf L}_{A}={\bf Q}_{L_{A}}$, ${\bf Z}_{A}={\bf R}_{L_{A}}{\bf Z}_{A}$, and ${\bf L}_{B}={\bf Q}_{L_{B}}$, ${\bf Z}_{B}={\bf R}_{L_{B}}{\bf Z}_{B}$. In such a way, ${\bf A}={\bf L}_{A}{\bf Z}_{A}$ and ${\bf B}={\bf L}_{B}{\bf Z}_{B}$ remain unchanged, while meeting the identifiability constraints ${\bf L}_{A}^{\intercal}{\bf L}_{A}={\bf I}_{R_{A}}$ and ${\bf L}_{B}^{\intercal}{\bf L}_{B}={\bf I}_{R_{B}}$. Column switching and sign switching was not an issue for the considered exercises, but could be readily resolved using order constraints during the mcmc or via ex post identification strategies. Finally, recalling Hoff (2015), ${\bf A}$, ${\bf B}$ and ${\bf\Sigma}_{A}$, ${\bf\Sigma}_{B}$ are only identified up to scaling, i.e., ${\bf B}\otimes{\bf A}=({\bf B}/c)\otimes(c{\bf A})$ and ${\bf\Sigma}_{B}\otimes{\bf\Sigma}_{A}=({\bf\Sigma}_{B}/c)\otimes(c{\bf\Sigma}_{A})$. To address this issue, we follow the recommendation within Appendix B of Hoff (2015) by rescaling ${\bf A}$, ${\bf B}$ and ${\bf\Sigma}_{A}$, ${\bf\Sigma}_{B}$ to maintain a constant relative magnitude among $||{\bf A}||^{2}$ and $||{\bf B}||^{2}$, and between $||{\bf\Sigma}_{A}||^{2}$ and $||{\bf\Sigma}_{B}||^{2}$, respectively.
5 5. Modeling the Criminal Activity Tensor in the City of Chicago
We conclude by illustrating the predictive performance and the inference potential of the Bayesian low-rank bilinear autoregressive model in Sections 34, with a focus on criminal activity data in the city of Chicago. These data are described in detail in Section 5.1. The results of the analysis under the proposed model can be found in Section 5.2.
5.1 5.1. Data
The three-way criminal activity tensor we rely on is available at https://www.chicagopolice.org/data-statistics/ and comprises time-stamped reports from the Chicago Police Department on $K=28$ crime categories within the $Q=22$ districts of the city. In our application, these data are aggregated on a monthly timescale from January 2001 until December 2019 in order to avoid structural breaks due to the onset of the covid-19 pandemic in the early 2020. Figure 1 provides a graphical representation of the dynamic log-counts of these reports for each crime category, aggregated over districts, whereas Figure 2 displays selected heatmaps of the number of crime reports across the $Q=22$ districts for specific categories, averaged over the time window analyzed. Consistent with the discussion in Section 2, both figures clarify that criminal activities are characterized by heterogeneous structures across time, crime categories and districts, which further showcase nuanced and complex interdependence patterns, thereby motivating the implementation of the model developed in Sections 34.
Notice that the $K=28$ crime categories we study in this application mainly follow the classification system employed by the Federal Bureau of Investigation (fbi), which aligns with the National Incident-Based Reporting System (nibrs). However, we observed that a nibrs category Miscellaneous Other often played an intriguing role in our results. To enhance the interpretability and specificity of our analysis, we therefore decided to redistribute crimes within this category based on the corresponding Illinois Uniform Crime Reporting (iucr) code. For example, we reallocated reports corresponding to iucr code 2092 – solicit narcotics on public way to the fbi category Drug Abuse (18), which therefore represents a broader, narcotics-related category. Furthermore, certain other iucr codes previously part of fbi category Miscellaneous Other have been bundled into new, separate, crime categories. An example for this are several codes related to threat and harassment, for instance 2820 - telephone threat or 2826 - harassment by electronic means. We reclassified these codes into an explicit Threat & Harassment category. This refinement of the crime classification allows us to capture the nuances of different crime types more accurately and facilitates a more in depth analysis.
Figure 1: For each crime category, log-transformed counts of police reports by month, aggregated across districts. The time window spans from January 2001 to December 2019.
Figure 2: Heatmaps of reports’ counts for selected crime categories in each district of Chicago, averaged over time. Data are shown on a logarithmic scale.
Before discussing the results of our analysis, it is also worth to mention that any study based on police reports provides only a partial understanding of the underlying criminal activity patterns due to several factors. First, police reports inherently represent the information on criminal activity filtered through the lens of law enforcement. As such, it is important to acknowledge that the reported incidents may not perfectly align with the actual criminal activity across the city. Second, the spatial dimension of these reports introduces complexities, as different neighborhoods may have varying levels of trust in the police, leading to differences in reporting patterns. Finally, the type of crime being reported and the temporal aspect can also impact the results, further highlighting the need for caution when interpreting the findings. Nonetheless, despite these limitations, analyzing police reports remains the standard approach in state-of-the-art literature, and is commonly used for forecasting and utilized by the law enforcement itself. It provides valuable insights into crime patterns, police activity, and the structures that emerge from the interaction among the two. That is, the results may be influenced by both the system of policing and the actual criminal activity, making it challenging to disentangle the two. Nonetheless, the interpretable model-based perspective we propose can uncover hidden patterns within the data produced by law enforcement, thereby providing police with more refined methods to examine and leverage information from these reports.
5.2 5.2. Results
The presentation of the results of the analysis under the proposed model is divided in four parts. In Section 5.2.1, we first compare the forecasting accuracy of model (1)–(2) with those achieved by a full-rank bilinear autoregressive formulation and a black-box predictive strategy relying on random forests. Besides clarifying that the proposed parsimonious representation is competitive with strategies specifically designed for forecasting, this assessment provides reassurance on the fact that inference on the interpretable dependence structures encoded in $\bm{\Sigma}_{A}$, $\bm{\Sigma}_{B}$, ${\bf A}$ and ${\bf B}$ relies on a realistic characterization of the observed criminal activity tensor. The results of inference on the posterior point estimates for these quantities are discussed in Sections 5.2.2 and 5.2.3, respectively. Finally, in Section 5.2.4 we interpret the findings arising from such inferences in the light of the general qualitative and quantitative evidence on criminal activities presented within Section 2.
All the results presented are based on $3{,}000$ posterior samples for the quantities of interest produced by Algorithm 1, following an initial burn-in period of $1{,}000$ iterations. Convergence and mixing of such an algorithm posed no major challenges. To select the ranks $R_{A}$ and $R_{B}$, we conducted an exhaustive enumeration of all possible values for the pair $(R_{A},R_{B})$, evaluating these combinations in terms of out-of-sample predictive accuracy, similarly to the assessments conducted in Section 5.2.1. The results of this analysis suggest $R_{A}=16$ and $R_{B}=8$. This is the lowest rank specification for the proposed model (1)–(2) that yields an out-of-sample predictive accuracy within $1\%$ of the one achieved by a fully-rank bilinear vector autoregressive formulation.
5.2.1 5.2.1. Predictive performance and benchmarking
Before discussing the results of inference under model (1)–(2), let us first assess its forecasting performance and compare it against alternative predictive strategies. In presenting the results of such an analysis, it is important to emphasize that while a significant portion of the empirical literature on crime data analysis is primarily designed for purely-predictive purposes, the model developed in Sections 34 aims instead at providing interpretable inference on the core dependence structures within the observed criminal activity tensor. As such, the overarching goal of our predictive assessment and comparisons is not to demonstrate that the proposed model systematically outperforms all potential competitors, including black-box ones, in out-of-sample forecasting. Rather, we aim to provide empirical evidence on the fact that carefully-designed model-based representations, such as the one we developed, can substantially enhance interpretability and inference potentials, while preserving a predictive accuracy that is competitive with those achieved by more sophisticated strategies, including highly-parameterized formulations and black-box solutions. Such an assessment also serves as a sanity check to quantify whether the proposed model yields a realistic representation of the generative mechanisms underlying the data analyzed, thereby reassuring on the reliability of the inference results in Sections 5.2.2 and 5.2.3.
mse mae corr
[…]
Full-rank bilinear autoregression 0.380 0.279 0.735
Random forests 0.367 0.272 0.755
[…]
Table 1: Forecasting performance of the proposed model (low-rank bilinear autoregression) and relevant competitors (full-rank bilinear autoregression, random forests) under different metrics: mse (mean squared error), mae (mean absolute error) and corr (correlation between predictions and actual observed values). Results are averaged across 24 replicated one-step-ahead forecasting exercises. The proposed low-rank bilinear autoregressive model employs $R_{A}=16$, $R_{B}=8$.
Consistent with the above goals, the predictive assessment is conducted by dividing the data into a training set and a test set. The training sets cover data ${\bf Y}_{1},\ldots,{\bf Y}_{t_{\mbox{\tiny train}}}$ for 204 consecutive months. Leveraging these training data we fit the competing methods under analysis, including the proposed model, and then produce one-step-ahead forecasts for the reported criminal activities at the time $t_{\mbox{\tiny train}}+1$ for all the 616 combinations of districts and crime categories. The resulting forecasts are finally compared to the actually observed values in ${\bf Y}_{t_{\mbox{\tiny train}}+1}$ using three different accuracy measures, namely the mean squared error (mse), the mean absolute deviation (mae) and the correlation (corr) between the forecasted values and actual observed ones in ${\bf Y}_{t_{\mbox{\tiny train}}+1}$. To provide a comprehensive assessment, these measures are further averaged over 24 replications of such a predictive analysis, each obtained by shifting the training window of 204 observations forward by one month, and forecasting the criminal activities in the subsequent month. As a result, predictive performance is quantified on a total of $616\cdot 24=14{,}784$ point forecasts.
Table 1 summarizes the outcome of the aforementioned predictive study for the proposed model (low-rank bilinear autoregression with $R_{A}=16$, $R_{B}=8$) and two relevant competing methods. The first one (full-rank bilinear autoregression) still relies on the bilinear representation as in (1), but does not include the low-rank structure in (2). The second (random forests (Breiman, 2001)) provides instead a relevant example of a routinely-implemented black-box strategy specifically designed for predictive purposes. As such, one would expect that this latter strategy would substantially improve the forecasting accuracy of the low-rank and full-rank bilinear autoregressive models. As illustrated in Table 1, this is not the case in such a context. In fact, both model-based solutions achieve a predictive accuracy competitive with random forests, with the low-rank formulation (1)–(2) providing essentially the same performance as its full-rank counterpart. These results clarify that interpretable model-based inference can be accomplished at negligible costs in predictive accuracy, and provides reassurance that the results presented within Sections 5.2.25.2.3 are reliable and not based on an unrealistic low-rank characterization of the data-generating mechanism.
5.2.2 5.2.2. Inference on intra-temporal dependence structures
Besides achieving satisfactory predictive accuracy, a distinctive feature of the model developed in Section 34 lies in its ability to disentangle complex dependencies in the criminal activity tensor, while allowing interpretable inference on such quantities. To showcase such potentials, let us first focus on the intra-temporal dependence structures among crime categories and between districts encoded in the posteriors of $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$. Figure 3 provides a graphical representation of these estimated dependences with a focus on the posterior mean of the partial correlation networks defined as $-\mbox{diag}(\bm{\Sigma}_{A}^{-1})^{-1/2}\bm{\Sigma}_{A}^{-1}\mbox{diag}(\bm{\Sigma}_{A}^{-1})^{-1/2}$ and $-\mbox{diag}(\bm{\Sigma}_{B}^{-1})^{-1/2}\bm{\Sigma}_{B}^{-1}\mbox{diag}(\bm{\Sigma}_{B}^{-1})^{-1/2}$ (e.g., Lauritzen, 1996). Such posterior means can be obtained by evaluating these formulas at the posterior samples of $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$ produced by Algorithm 1 and then averaging across samples. Crucially, these measures provide an informative reconstruction of the graphical dependencies between crime categories and across the different districts, which quantify the intra-temporal pairwise correlations among any two generic nodes in the network conditionally on the effect of all the others, and net of the inter-temporal dependencies captured by the conditional mean in (1)–(2).
Figure 3: Estimated intra-temporal partial correlation networks between crime categories (left) and among districts (right). The colors in each network indicate clusters of nodes with high within-group absolute partial correlations and relatively-low across-group dependencies (the clusters are estimated via community detection algorithms (Blondel et al., 2008)). Node positions are obtained via force-direct placement methods (Fruchterman and Reingold, 1991) positioning closer in the space nodes with higher absolute partial correlations.
Figure 4: Left: Districts of Chicago colored according to the communities estimated from the partial correlation network in the left panel of Figure 3. Right: Plot of the log-transformed distances between pairs of district centroids versus the posterior mean of the partial correlation among such district pairs (loess estimate and uncertainty bands for the relation between these two quantities are also provided).
Figure 3 represents graphically the aforementioned networks with nodes corresponding to crime categories and districts, respectively, and weighted edges denoting the posterior mean of the partial correlation among any pair of nodes. To facilitate interpretation and uncover relevant structures, the positions of the nodes within Figure 3 are obtained via force-direct placement methods (Fruchterman and Reingold, 1991), so that nodes with higher absolute partial correlations are displayed closer in the space. This graphical setting is also combined with the output of a community detection algorithm (Blondel et al., 2008) that unveils clusters of nodes having high within-group absolute partial correlations and relatively-low across-group dependencies. As shown in the network among the different crime categories, such a perspective allows us to uncover yet-unexplored structured modular intra-temporal dependencies between types of crimes. More specifically, the first group (green nodes) comprises crimes such as simple and aggravated battery and assault, which are physical crimes directly targeting the body. The second (purple nodes) includes, instead, crimes related to narcotics, gambling, weapons violations, and liquor license-related offenses. This clustering suggests a potential association with organized crime or coordinated gang activities in the city. The third group (blue nodes) consists mainly of property crimes such as burglary, robbery, theft, vandalism, and trespassing. These types of crime are often interconnected for various reasons. For example, the evidence suggests that the presence of vandalism, like graffiti, can significantly increase the propensity of people to engage in property crimes, such as stealing (Keizer et al., 2008). At the same time, it is worth noting that the connection could be directly related to legal definitions. To give an instance, trespassing evolves into burglary when an individual trespasses with the intent of committing a crime during that trespass (i.e., the difference pertains to whether the crime in question was attempted or actually committed). Finally, the last group (red nodes) captures crimes which exhibit sparse correlations, even within the community itself. This cluster encompasses crimes such as arson, homicide or kidnapping, which are offenses that tend to occur independently of other crimes. A partial exception in such a group is provided by prostitution, which appears in Figure 3 relatively close to the purple nodes, thereby suggesting that this category might have also direct connections with organized crime or coordinated gang activities. See Section 5.2.4 for further discussion on these findings in the light of the qualitative and quantitative criminology theories presented in Section 2.
Similarly-interesting results can be obtained by examining the network structure of the partial correlations among the districts displayed in the right panel of Figure 3. The community detection algorithm (Blondel et al., 2008) applied to such a network reveals in this case two distinct groups, which fully align with the north-south division of the city of Chicago. This geographical separation becomes even more evident when visualizing the estimated group structure on a map of the city, as demonstrated in the left panel of Figure 4. The connection of the partial correlations to geographical distance is further underlined in the right panel of Figure 4, which highlights a negative relationship between the pairwise distances among the districts and the corresponding partial correlation. As discussed in Section 5.2.4 these findings contribute to the ongoing studies on criminal activities within urban environments and provide also useful resources to quantify potential spillover effects across the districts. Remarkably, the proposed formulation identifies these spatial dependence structures without including any explicit geographical information within the model. Although it would be interesting to extend the proposed framework to explicitly incorporate this additional source information, such an extension is beyond the scope of the current article and might even undermine the flexibility of the model. In fact, the spatial dependencies that we learn might arise from the direct effect of underlying latent factors such as localized activities of criminal organizations or distinctive policing strategies, which are often difficult to include in the model only through geographical information. In fact, avoiding such constraints allows us to learn in the left panel of Figure 4 stronger dependencies among southern districts relative to those inferred for northern ones. This result partially departs from pure geographical proximity and provides important quantitative evidence on the presence of organized crime and structured gangs activities in the southern area of Chicago (e.g., Aspholm, 2019).
5.2.3 5.2.3. Inference on inter-temporal dependence structures
Figure 5 displays the heatmaps of the posterior means for the elements in the matrices ${\bf A}$ and ${\bf B}$. Recalling Section 3.1, these quantities provide insights into the average inter-temporal dependence among the $K$ crime categories within districts, and between the $Q$ districts for each crime category, respectively. As such, although these estimates may not directly correspond to causally-interpretable effects, the results in Figure 5 are still able to unveil relevant association patterns. First, the dominant feature of both matrices is the main diagonal, indicating a moderately-high inter-temporal persistence within reports by crime categories and districts, respectively. In the posterior mean of the matrix ${\bf A}$ (see left panel of Figure 5) certain crimes, such as arson and homicide, occur infrequently and in relative isolation, leading to exceptions in this overall pattern. Additionally, the occurrence of specific crime categories, such as simple battery, vandalism, and theft-related crimes, appears to have a more remarkable association with the potential outbreak of specific types of crimes in the subsequent month. Such a finding provides important guidelines in the design of intervention policies guided by potential spillover effects.
Similarly interesting results are provided by the posterior mean of the matrix ${\bf B}$ in the right panel of Figure 5. Besides the direct inter-temporal persistence highlighted in the main diagonal, the estimated ${\bf B}$ clarifies that such a persistence acts also at a group level, especially for clusters of southern districts whose crime activities appear to be associated with those at subsequent times in other districts from the same group. This result is interestingly in line with the spatial intra-temporal dependencies discussed in Section 5.2.2 and further support the idea that these patterns might arise from the presence of organized crime and coordinated gangs activities in the southern area of the city (e.g., Aspholm, 2019).
Figure 5: Heatmaps of the posterior mean for the matrices ${\bf A}$ (left) and ${\bf B}$ (right) characterizing the inter-temporal dependencies under model (1)–(2).
5.2.4 5.2.4. Discussion of the results
The results in Sections 5.2.25.2.3 provide novel empirical evidence in the context of criminological analyses on the mechanisms and determinants driving concentration of specific crime types in time and space (e.g., Brantingham and Brantingham, 2016, 2017; Cohen and Felson, 1979). Social, economic, and even structural characteristics of urban environments influence criminal behavior (Brantingham and Brantingham, 1995), leading to specific inter and intra-temporal dependencies within both crime categories and spatial districts that also exhibit peculiar modular patterns (see Figure 3 and Figure 5). As such, our results further underscore the importance of maintaining a systemic perspective on crime phenomena, accounting for all the three dimensions of the observed criminal activity tensor, while inferring interpretable dependencies within and across such dimensions. For example, the spatial effects we unveil in Figure 3 and Figure 5 highlight an heterogeneous distribution of crime generators and crime attractors within the urban environment (Brantingham and Brantingham, 1995; Weisburd, 2015), which, in turn, do not facilitate the same crime categories across the whole city. An interesting case is provided by the Chicago Loop located in the 1st district. This area serves as a financial, cultural, and commercial hub of the city, thereby increasing the number of potential targets, especially in relation to property crime. The right panels of Figure 3 and Figure 5 show that, although the 1st district is located in the northern group, it still provides a crime generator area that showcases non-negligible intra- and inter-temporal associations with southern districts. Recalling, e.g., Aspholm (2019), these southern districts tend to act, instead, as crime attractors due to reduced interventions and structural organizational issues that create opportunities for organized criminal activities, and hinder residents from exerting informal social control (Sampson and Groves, 1989; Shaw and McKay, 1942). North Lawndale, situated in the 10th district (southern group in Figure 3 and Figure 4), exemplifies this dynamic. As one of the most impoverished and socially disorganized neighborhoods in Chicago, it experiences high rates of homicides and gun violence. These criminogenic disparities across the different areas of Chicago seem to emerge in our findings, which highlight clear modular separations in the intra- and inter-temporal associations within and across northern and southern Chicago districts, with the latter exhibiting stronger within-group dependencies.
The results within Figure 4 further underscore the importance of the geographic proximity and potential spillover effects. According to our findings, there is a negative relationship between the distance separating pairs of districts and the corresponding partial correlation. One possible explanation involves mechanisms that combine residential location and social interactions (Zenou, 2003), highlighting the role of behavior in disadvantaged environments (Glaeser et al., 1996). For instance, evidence suggests that offenders are inclined to exploit criminal opportunities close to their residential areas (e.g., Bernasco, 2010; Bernasco and Block, 2009; Bernasco et al., 2014; Hipp, 2016; Menting et al., 2016; Townsley et al., 2016). Consequently, the concentration of potential offenders within a neighborhood can trigger a contagion effect, propagating criminal behavior and reinforcing negative social dynamics (Becker and Murphy, 2009).
Moreover, increased criminal concentration intensifies the competition among offenders, reducing the expected gain of criminal activity (Becker, 1968) and encouraging repeat offenses (Zenou, 2003). This is also consistent with research on repeat victimization and temporal clustering, which indicates that the occurrence of an initial crime increases the probability of subsequent offenses in the same location. Offenders tend to exploit previously-acquired knowledge about susceptible targets and the absence of effective interventions (e.g., Johnson and Bowers, 2004; Townsley et al., 2003). This persistence of crime is also evident in our findings (see Figure 5), although there are differences across crime categories. Specifically, while the main diagonal of the left matrix within Figure 5 shows a clear temporal persistence for property crimes such as theft, robbery, and burglary, as well as for organized crime-related offenses like narcotics, prostitution, and gambling, this pattern is less pronounced for the violent offenses such as sexual-related crimes, homicides, and manslaughter, and for the property crimes like arson and embezzlement. These inferred heterogeneities in crime persistence by offense type provide fundamental empirical evidence to the literature on such phenomena (Chainey, 2021; Haberman and Ratcliffe, 2012; Johnson and Bowers, 2004; Johnson et al., 2009).
6 6. Conclusion
In this article, we proposed a low-rank bilinear autoregressive model to uncover interpretable dependence structure within and across the three dimensions (time, space, crime categories) of observed criminal activity tensors, with a specific focus on report data from the city Chicago. Besides disentangling and facilitating inference on complex interdependence patterns over several dimensions, the proposed formulation crucially achieves also a predictive accuracy that is competitive with black-box strategies, thereby providing a realistic, yet interpretable, model-based representation of crime phenomena over space, time and different categories. The results in the application confirm these advantages, and further stress the importance of developing joint models that maintain a systemic perspective on crime phenomena by formally accounting for dependencies across and within all the tree dimensions of a generic criminal activity tensor. This perspective not only contributes to enlarging available knowledge on the mechanisms and determinants of crime, but also facilitates the design and evaluation of improved policies and intervention strategies.
The above advancements motivate also several directions for future research. First, extending the model to incorporate nonlinearities could provide further insights into the complex dynamics of criminal activities that might also yield further gains in predictive accuracy. Relatedly, another interesting direction for future research would be to explicitly incorporate additional structural information, such as knowledge on activity patterns of organized criminal groups or spatial structure, into the modeling framework. The latter could be achieved by introducing external covariates or by incorporating latent spatio-temporal dependencies into the model parameters, leveraging, e.g., results in Hsu et al. (2021). Finally, given the typical dominance of the main diagonal in the estimated matrices ${\bf A}$ and ${\bf B}$, future research could consider other alternatives to the factorization in (2), such as, for example ${\bf A}={\bf D}_{A}+{\bf L}_{A}{\bf Z}_{A}$ and ${\bf B}={\bf D}_{B}+{\bf L}_{B}{\bf Z}_{B}$, respectively, with ${\bf D}_{A}$ and ${\bf D}_{B}$ diagonal. In terms of applications, it would be interesting to develop quantitative strategies for designing and assessing intervention policies guided by the output of the proposed model, especially with reference to the magnitude and structure of potential spillover effects. Expanding the analysis to other crime datasets from different cities and regions would also be important to obtain a more comprehensive perspective on criminal activities in urban environments beyond the city of Chicago. Moreover, the proposed model and analyses can also be applied beyond crime, whenever data are available in the form of activities/choices/behaviors/events over space and time with respect to a number of categories. An example could be employment data in urban environments over time and districts, with respect to different job categories.
7 Acknowledgments
Carlos Díaz gratefully acknowledges financial support from the National Agency of Research and Development’s (ANID) Millennium Science Initiative Program (NCS2024 058). Daniele Durante is funded by the European Union (ERC, NEMESIS, project number: 101116718). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.